OpenStudy (anonymous):

Find the x intercepts for the graph if the equation y=x^2+2x-3

OpenStudy (anonymous):

replace y with zero, then solve for x

OpenStudy (anonymous):

x^2 + 2x - 3 = 0 (x + 3) (x - 1) = 0 x = -3, 1

OpenStudy (anonymous):

thats what i got thx